OpenStudy (anonymous):

Can someone please work out this equation step by step for me? I keep getting my math wrong somewhere. The correct answer is -2.

OpenStudy (anonymous):

\[\sqrt{3x+10} = \sqrt{x+11} - 1\]

OpenStudy (anonymous):

Let me try it really quickly

OpenStudy (anonymous):

Okay! Thank you!

OpenStudy (anonymous):

Ok so I got negative 2 but I got another answer with it....x=5

OpenStudy (anonymous):

Anyone want to confirm that with me?

OpenStudy (anonymous):

I have no idea what the heck I did wrong. Let me go back over my math.

OpenStudy (anonymous):

Well I mean I was asking more so if they gave another answer since like I said I got another answer with the -2

OpenStudy (anonymous):

No, it just says -2 is the only answer.

OpenStudy (anonymous):

Let me recalculate it a different way to see what I get

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

Hm....I seem to be having a brain fart since I can't seem to get a stable answer....but I'm still going to try it

OpenStudy (anonymous):

Haha okay. Yeah my brain doesn't want to work today...

OpenStudy (anonymous):

I dont know why I'm having so much trouble with this....

OpenStudy (anonymous):

Yeah me either.

OpenStudy (anonymous):

Yeah sorry for some reason I cant figure it out

OpenStudy (anonymous):

It's okay. No big deal.